# Importance Ranking in Complex Networks via Influence-aware Causal Node Embedding

### v2

@@ 5.6.3 Ablation study for the causal ranking loss
Remark: Our work follows a similar paradigm to the work of AGNN [xiong2024vital] and STGCN [huang2025sparse]. Both frameworks achieve superior node importance ranking by integrating graph neural networks with autoencoder-based representation learning and transfer learning from synthetic to real-world networks. The novelty of our model compared to these two is that we introduce an influence-aware causal node embedding mechanism to ensure that the learned embeddings are causally related to node influence. We further propose a unified objective function that integrates a causal reconstruction loss and a causal ranking loss with a regularization term. This framework synergistically optimizes both feature representation and ranking, which is validated by ablation studies to enhance ranking performance and generalization capability.
@@ 5.8 Parameter sensitivity analysis
To investigate the sensitivity to the balanced parameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$, we conduct a parameter sensitivity analysis of the models trained on the BA network, evaluating their performance on various real-world networks.
Let $\lambda_{2}=\lambda_{3}=1$, the Kendall’s coefficient with different values of $\lambda_{1}$ are shown in Fig. 3(a). Similarly, setting the values of the other two parameters to be the same, the results with difference values of $\lambda_{2}$ and $\lambda_{3}$ are displayed in Fig. 3(b) and Fig. 3(c) respectively. As shown in the figures, ICAN is sensitive to the parameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$. Specifically, for $\lambda_{1}$, the optimal value range is $[0.3,0.5]$ for Karate, and $[0.5,1.5]$ for the other networks. As for $\lambda_{2}$, the range of $[0.1,1.5]$ is optimal for the Vidal, as well as the Email-univ and Email-dnc networks, $[0.1,0.3)$ for Jazz, $[1,1.5)$ for USAir and Karate. Regarding $\lambda_{3}$, the optimal value range is $(0.3,0.5]$ for Jazz, $[1,1.2)$ for USAir,and $[0.5,1.5]$ for the other networks. Based on these findings, in the implementation of ICAN, the parameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ can be set to (0.5, 1, 1.5) for the Karate network, (1, 0.1, 0.5) for the Jazz network, (0.5, 1, 1) for the Email-univ network, (1, 1, 1) for the other networks, respectively.
@@ 6 Discussion
In this section, we discuss the implications of the proposed model from both network science and engineering perspectives. First, we examine the degree distribution of the most influential nodes identified by the model, providing insights into the structural characteristics that underlie node importance. We then explore the model’s behavior and applicability in social networks with node attributes, highlighting its potential in more complex, attribute-rich environments.
@@ 6.1 Degree distribution of Top 10% Influential Nodes
A clear methodological contrast emerges. As expected for the degree–centrality-based approach, the results for DC in Figs. 4(b)–6(b) exhibit a strong positive correlation: nodes with the highest influence scores coincide with those of the largest degrees, concentrating red points at the high-degree end of the spectrum. In contrast, ICAN and CGNN reveal markedly different patterns. High-scoring nodes appear across a broader degree range, with a noticeable concentration in the moderate-degree region. This suggests that these methods attribute high influence to nodes that are not necessarily the network’s primary hubs.
@@
6.2 Case study for social network with attributes
In this case study, we show that ICAN can be applied to a social network with attribute information. The feature matrix can be constructed from node attributes. Cora [cabanes2012cora] is a citation network where nodes represent scientific publications and edges represent directed citation links between them. Each node feature in the Cora dataset is a 1,433-dimensional binary bag-of-words vector representing the presence or absence of specific terms in the paper’s content. Our model is trained on BA synthetic networks (using Node2Vec for feature generation) and evaluated on the attributed Cora real-world network. The experimental results, as shown in Fig. 7, demonstrate that our method achieves superior ranking performance compared to all baseline approaches. The results demonstrate that ICAN can be applied to social networks with attribute information, indicating its potential for broader applications in social network analysis. However, in such attribute-rich networks, inconsistencies may arise between topological structures and node attributes [ma2025node]. Therefore, future work will focus on developing more effective embedding strategies that can integrate both sources of information in a coherent manner.
@@ 7 Conclusions
Quantifying node importance is a challenging task, particularly in non-cooperative scenarios where the target network structure is not accessible in advance due to privacy or security constraints. This paper proposes a node importance ranking method for complex networks based on an influence-aware causal autoencoder model. The proposed method effectively captures the network-invariant causal relationships between node embeddings and influence scores. This design allows ICAN to learn robust, low-dimensional node representations solely from synthetic networks, which can then be applied to node-ranking tasks on various real-world networks. We conduct comprehensive experiments by training ICAN on different types of synthetic networks and evaluating it on diverse benchmark networks, comparing its performance with eleven representative node-ranking methods. The results confirm that ICAN achieves state-of-the-art performance, excelling in both ranking accuracy and cross-network generalization. The results of the ablation studies confirm that both the influence-aware causal node embedding mechanism and the feature–task co-optimization strategy—based on the jointly defined causal ranking loss and causal reconstruction loss—play crucial roles in ensuring that the learned representations generalize effectively across diverse target graphs and are better suited for downstream ranking tasks, thereby substantially enhancing the model’s overall performance. Future work will focus on establishing a theoretical framework for selecting or generating suitable training networks for a given network engineering task. In addition, subsequent research will undertake a deeper investigation of ICAN in practical network engineering contexts, such as developing cost-aware strategies for network dismantling and designing embedding methods that effectively integrate topological and semantic information in social networks.

### v3

@@ 5.6.3 Ablation study for the causal ranking loss
Remark: Our work follows a similar paradigm to the work of AGNN [69] and STGCN [29]. Both frameworks achieve superior node importance ranking by integrating graph neural networks with autoencoder-based representation learning and transfer learning from synthetic to real-world networks. The novelty of our model compared to these two is that we introduce an influence-aware causal node embedding mechanism to ensure that the learned embeddings are causally related to node influence. We further propose a unified objective function that integrates a causal reconstruction loss and a causal ranking loss with a regularization term. This framework synergistically optimizes both feature representation and ranking, which is validated by ablation studies to enhance ranking performance and generalization capability.
@@ 5.8 Parameter sensitivity analysis
To investigate the impact of the balanced parameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$, we conduct a parameter sensitivity analysis of the models trained on the BA network, evaluating their performance on eight real-world networks. Let $\lambda_{2}=\lambda_{3}=1$, the Kendall’s coefficients with different values of $\lambda_{1}$ are shown in Fig. 3(a). Similarly, setting the values of the other two parameters to be the same, the results with different values of $\lambda_{2}$ and $\lambda_{3}$ are displayed in Fig. 3(b) and Fig. 3(c) respectively. The results demonstrate that while the model responds to parameter variations, it exhibits significant robustness and consistency across the majority of datasets. Specifically, for $\lambda_{1}$, the performance remains stable and optimal within the range $(0.5,1.2)$ for all networks. Similarly, for $\lambda_{2}$ and $\lambda_{3}$, most networks (7 out of 8) achieve optimal or near-optimal performance when the parameters are set around $1$. The Jazz network is the only exception, requiring lower values ($\lambda_{2}\approx 0.1,\lambda_{3}\approx 0.5$) to reach its peak, likely due to its distinct topological structure compared to the other networks. In practical scenarios without ground truth labels, users need a reliable default setting. Based on the extensive overlap of optimal regions across the majority of datasets, we empirically recommend a unified setting of $(\lambda_{1},\lambda_{2},\lambda_{3})=(1,1,1)$. This configuration effectively balances the model components and ensures robust performance for most real-world networks, eliminating the need for extensive parameter tuning in most cases.
To investigate the effect of different parameterizations for synthetic training networks, we generated BA networks with varying number of nodes (1,000, 2,000, and 3,000) and average degrees (2, 4, and 8). Their performance is evaluated on six real-world datasets, as shown in Fig. 4. The results indicate that the model yields the best performance with 1,000 nodes and an average degree of 4. Based on these findings, we recommend this parameter as the default training setting for practical applications.
Figure 4: Performance comparison of different training network parameters (numbers of nodes and average degrees) on six real-world networks.
@@ 6 Discussion
In this section, we discuss the implications of the proposed model from both network science and engineering perspectives. First, we examine the degree distribution of the most influential nodes identified by the model, providing insights into the structural characteristics that underlie node importance. We then explore the model’s behavior and applicability in real-world networks with structural descriptors as node features and social networks with node attributes, highlighting its potential in more complex, attribute-rich environments.
@@ 6.1 Degree distribution of top 10% influential nodes
A clear methodological contrast emerges. As expected for the degree–centrality-based approach, the results for DC in Figs. 5(b)–7(b) exhibit a strong positive correlation: nodes with the highest influence scores coincide with those of the largest degrees, concentrating red points at the high-degree end of the spectrum. In contrast, ICAN and CGNN reveal markedly different patterns. High-scoring nodes appear across a broader degree range, with a noticeable concentration in the moderate-degree region. This suggests that these methods attribute high influence to nodes that are not necessarily the network’s primary hubs.
@@
6.2 Case study
Although ICAN utilizes node2vec embeddings as the feature matrix in our primary experiments, it is capable of flexibly accommodating diverse feature inputs. Specifically, for non-attributed networks, it can leverage explicit topological descriptors, while for attributed networks, it can directly utilize the raw attribute matrix. To substantiate this flexibility, we conduct two supplementary experiments in this subsection.
6.2.1 Experiments with structural descriptors as node features
In this experiment, we select six classical node centrality measures as feature vectors to represent the characteristics and structural properties of nodes, including degree centrality, betweenness centrality, closeness centrality, clustering coefficient, eigenvector centrality, and load centrality. To reduce the risk of overfitting, feature values for each node are normalized using min-max scaling so that they fall within the interval (1,1)(-1,1). The resulting normalized centrality vectors form the feature matrix fed into ICAN. We train ICAN on BA synthetic network and evaluate its generalization performance on eight real-world networks. As shown in Table VII, the variant using structural descriptors as node features (ICAN-S) achieves slightly lower average ranking performance than the node2vec-based version (ICAN). However, as reported in Table II, ICAN-S consistently achieves superior average ranking performance over all other baselines across datasets. These results indicate that ICAN can flexibly adapt to different types of input features while maintaining strong generalization and robust ranking performance.
TABLE VII: kendall’s $\tau$ of ICAN and ICAN-S on eight real-world networks.
Methods Karate Jazz Email-univ USAir Vidal Email-dnc Figeys Oz Average
ICAN 0.8090 0.8504 0.7625 0.6758 0.9524 0.5740 0.6708 0.8051 0.7625
ICAN-S 0.8033 0.8259 0.6598 0.6338 0.9524 0.5939 0.6391 0.7819 0.7363
6.2.2 Social network with attribute-enriched node features
In this case study, we show that ICAN can be applied to a social network with attribute information. The feature matrix can be constructed from node attributes. Cora [7] is a citation network where nodes represent scientific publications and edges represent directed citation links between them. Each node feature in the Cora dataset is a 1,433-dimensional binary bag-of-words vector representing the presence or absence of specific terms in the paper’s content. Our model is trained on BA synthetic networks (using node2vec for feature generation) and evaluated on the attributed Cora real-world network. The experimental results, as shown in Fig. 8, demonstrate that our method achieves superior ranking performance compared to all baseline approaches. The results demonstrate that ICAN can be applied to social networks with attribute information, indicating its potential for broader applications in social network analysis. However, in such attribute-rich networks, inconsistencies may arise between topological structures and node attributes [46]. Therefore, future work will focus on developing more effective embedding strategies that can integrate both sources of information in a coherent manner.
@@ 7 Conclusions
Quantifying node importance is a challenging task, particularly in non-cooperative scenarios where the target network structure is inaccessible in advance due to privacy or security constraints. This paper proposes a node importance ranking method for complex networks based on an influence-aware causal autoencoder model. The proposed method effectively captures the network-invariant causal relationships between node embeddings and influence scores. This design allows ICAN to learn robust, low-dimensional node representations solely from synthetic networks, which can then be applied to node-ranking tasks on various real-world networks. We conduct comprehensive experiments by training ICAN on different types of synthetic networks and evaluating it on diverse benchmark networks, comparing its performance with eleven representative node-ranking methods. The results confirm that ICAN achieves state-of-the-art performance, excelling in both ranking accuracy and cross-network generalization. The results of the ablation studies confirm that both the influence-aware causal node embedding mechanism and the feature–task co-optimization strategy—based on the jointly defined causal ranking loss and causal reconstruction loss—play crucial roles in ensuring that the learned representations generalize effectively across diverse target graphs and are better suited for downstream ranking tasks, thereby substantially enhancing the model’s overall performance. Future work will focus on establishing a theoretical framework for selecting or generating suitable training networks for a given network engineering task. In addition, subsequent research will undertake a deeper investigation of ICAN in practical network engineering contexts, such as developing cost-aware strategies for network dismantling and designing embedding methods that effectively integrate topological and semantic information in social networks.